# Embodied Image Captioning: Self-supervised Learning Agents for Spatially Coherent Image Descriptions

### v1

@@ Abstract
{strip}[Uncaptioned image]
Figure 1: An agent equipped with an instance segmentation and captioning model navigates the environment. Off-the-shelf captioners predict partially wrong and inconsistent captions across different views of the same object: CoCa [1], BLIP2 [2], Florence2 [3], and ChatGPT [4].
@@ I Introduction
Visual understanding is a fundamental skill to recognize and describe objects in a scene, enabling robots to navigate and interact with the environment [5]. Despite advances in image captioning [1, 6, 7, 8, 3], a captioner deployed on an autonomous agent often generates wrong or inconsistent descriptions across different views of the same object, especially in the case of occlusions or challenging viewing directions (see Fig. 1).
Previous works tackled the problem of inconsistent captions from three research directions: agent navigation [9, 10], noisy captions [6, 11, 12], and caption comprehensiveness [13, 14]. Navigation-based captioning methods combine the agent observations at different times during navigation [10]. However, when the agent moves in complex environments looking at the same object, the generated descriptions can be highly inconsistent. Methods that generate the most appropriate caption for an image either use the alignment between vision and language information to select the description from a set of captions [12], or prompt a Large Language Model (LLM) to summarize the captions [13]. Mentioned methods fail in case of inaccurate descriptions that might be selected or propagated to the captions summary.
Compared to previous works, we tackle both the noise presence and the caption coherence as challenges inherently present when a robot navigates a complex indoor environments. To this end, we propose a three-phase framework that combines embodied perception with a self-supervised strategy to retrain the captioner enhancing the performance and increasing the consistency of generated descriptions.
In the first phase, the agent equipped with an object detector and a captioner navigates the environment, collecting observed RGB-D images and corresponding object captions. The agent also creates a point cloud representation of the scene and clusters the object instances in 3D, given the output of the object detector. In the second phase, for each 3D segmented object instance, we prompt an LLM using the corresponding captions, their frequency and in-context learning to describe each object in the environment, hence filtering out potentially wrong captions and generating a comprehensive pseudo-label. In the third phase, we specialize the captioner through contrastive learning, maximizing the similarity of representation for different views of the same object while minimizing the similarity with other objects’ representation. The framework design is modular to allow the integration of past and future approaches for exploration, pseudo-caption generation, and environment specialization, fostering fair comparisons at each phase, and offering a foundation to improve caption consistency in challenging environments.
@@
  • A fine-tuning strategy that uses contrastive learning to improve caption consistency for different views of the same objects.
@@ II Related Works
Image captioning. Captioners adapt existing models by using additional inputs such as questions about the image, or bounding boxes  [8, 2, 15, 16, 17, 18, 19, 20, 21] or by improving the representations using contrastive learning [1, 22, 7, 8, 3, 6]. To fully exploit the generalization capabilities of foundation models, the majority of methods adjusts the representation of vision-language models fine-tuning tokens or transformer layers, reducing the gap between the pre-trained visual and text representations [2]. For example, X-InstructBLIP and BLIP2 bridge the representation between the image encoder and the text decoder by learning a set of queries to combine visual features with textual features using a lightweight transformer [8, 2]. Other methods like SCA and CAT adapt the SAM architecture [23] through visual chain-of-thoughts [17] or additional cross-attention layers to combine the vision and text representation and predict regional captions [15]. A common approach among captioning methods is to learn the representation through contrastive learning i.e., maximizing the similarity among input modalities such as vision, language, and point cloud [1, 22, 7, 8, 3, 6]. The multimodal alignment of different object representation can be learned either during a pre-training stage enabling downstream tasks such as captioning or classification [7, 8, 3, 6], or during the training to ground the captions to the visual features [1, 22]. Despite the use of contrastive learning on input modalities, methods fail to provide consistent captions across different views of the same object, generating wrong descriptions in the case of challenging viewpoints. In this work, we apply contrastive learning to induce the model to learn close representations in the latent space the same object while maintaining a margin distance from the representations of other objects, using triplet loss [24].
@@
Embodied captioning. Embodied agents navigate the environment and collect different views of the same object to learn a 3D representation of the scene [9] or to generate a caption [10]. For example, CaBOT decouples the policy that predicts the action to take in the environment, from the captioning model that generates the object description based on the trajectory [10]. Despite generating a single caption for an object, CaBOT has prior knowledge of the best viewpoint in the scene and scenes contain few objects, limiting the application in case of challenging indoor environments like houses or office rooms and in case the annotation of the best viewpoint is not provided.
@@ III Method
We propose a three-phase pipeline to improve the captioning performance and the consistency of captions across different views of the same object (see Fig. 2). In the first phase, an agent explores an environment following a policy $\pi:S_{t}\rightarrow a_{t}$, where $S_{t}$ is the state and $a_{t}$ the action at time instant $t$, and collects RGB-D observations. The agent is equipped with an instance segmentation model $f_{d}$ and an image captioning model $f_{c}$. $f_{d}$ restricts the RGB observation $I_{t}\in\mathbb{R}^{W\times H\times 3}$ of the agent to detected object instances $o$, $f_{d}:I_{t}\rightarrow\{{l}^{o}_{t},b^{o}_{t},M^{o}_{t}\}$, where $l\in\mathbb{R}^{O}$ is the logits vector of the $O$ classes $f_{d}$ can detect, $b\in\mathbb{R}^{4}$ is the object bounding box, and $M\in[0,1]^{W\times H}$ is the object segmentation mask. $f_{c}$ generates the description of each detected object $f_{c}:I^{\prime o}_{t}\rightarrow c^{o}_{t}$ where $I^{\prime}\in\mathbb{R}^{W^{o}\times H^{o}\times 3}$ is the object crop and $c$ the caption. During exploration, the agent builds a semantic voxel map of the environment consisting of $V$ points, $v_{t}\in\mathbb{R}^{V\times 3}$, and clusters the voxels into object IDs based on the detection logits. In the second phase, an LLM predicts a unique and concise description (pseudo-caption $\tilde{c}$) of each clustered object based on all the corresponding captions collected during exploration. In the third phase, the captioner is fine-tuned using the pseudo-captions with the goal of improving the consistency of captions for the same object instance.
@@ III-A Navigation and point cloud segmentation
Learned policy (Caption-Look Around - CLA). The built voxel map can be used to guide an exploration policy based on maximizing the disagreement between captions of the same object (Fig. 2a). For this reason, we adapt Look-Around [25] policy to tackle the captioning task. The policy takes an input $S_{t}$ defined as:
@@
where the first channel of $D^{map}_{t}$ contains a disagreement map with canonical dimension $K\times K$, where $K=128$; the second channel contains the map of the explored environment with the agent’s position superimposed; finally $e_{t}$ contains the global orientation of the agent with respect to the map. For the disagreement map, in this work we use a per-pixel measure of inconsistency between captions, computed as follows: for each object, all captions are encoded via SBERT [26]. Then, the cosine distance among all caption embeddings belonging to the same object is used as a measure of disagreement, which is down-projected onto a 2D image (see Fig. 3d). The policy learns to predict a long-term goal $(x,y)$ in the environment through reinforcement learning using the total disagreement as a curiosity reward [25].
@@ III-B Language-Driven Consistent PSeudocaptioner
The exploration phase assigns a unique ID to the object instances in the built voxel map. However, different captions $c^{o}_{t}$ are still associated with each object instance, and hence they are not consistent across views. We ensure consistency at the label level using a fully self-supervised approach based on prompting an LLM to perform caption distillation. We call our method LD-CPS (Language-Driven Consistent PSeudo-captioner). For each object instance, LD-CPS generates the pseudo-caption $\tilde{c}$ from the list of predicted captions $c_{t}$ as follows: the captions are first pre-processed to remove all uninformative text such as “A picture of…” or “A photo of…” that could be the result of bias in the captioner training or due to the distribution shift between training set and simulated environment. Then, a prompt is devised that asks an LLM to combine the list of all captions, given their frequency i.e., how many times they occur for the same object (Fig. 2d), into a concise caption (Fig. 2e). The prompt also highlights to the LLM the fact that some captions might be inaccurate or wrong. In addition to considering caption frequency, we use in-context learning, by providing a list of examples of how to solve the task inside the prompt. Despite using an LLM to summarize the caption, the proposed approach is different from IC3 [13], since we use the frequency of the captions in the prompt, under the assumption that the noisy data are not the majority. In-context examples are also provided to improve the pseudo-caption distillation task. We show the full prompt in Sec. D of the Supplementary Material.
@@ III-C Learning captions consistency
where $\lambda_{tr}$ is the coefficient weighting the contribution of the triplet loss.
@@ IV-A Datasets
We validate our framework on two available datasets: Gibson [27] and Habitat-Matterport (HM3D) [28]. Gibson reflects the complexity of 572 real-world environments featuring object categories such as tables, sofas, plants, in different settings (rooms and object instances). To avoid using degraded object meshes, we consider a total of 194 scenes with a reconstruction quality between 3 and 5 (the maximum), and we randomly split the scenes into training set (134), validation set (28) and test set (32) keeping the initial distribution of scene quality across the splits. HM3D comprises 1,000 reconstructed indoor residential and commercial environments having higher quality than Gibson [28]. We use the 181 scenes from the Habitat Challenge, split into a training set (126), a validation set (27), and a test set (28). We split data based on the scenes to evaluate the generalization of the framework to environments unseen during training in both Gibson and HM3D.
@@ IV-B Implementation details
Navigation. We import the datasets scenes in the Habitat simulator [29] that allows agent to explore environments and to collect observations. We compare three exploration policies: random goals, frontier exploration and the learned policy (CLA) presented in Sec. III-A. The first two policies use the same path planner that first converts the explored map into a visibility graph (nodes represent feasible goals and edges the straight trajectory without obstacles) via skeletonization, then computes the optimal trajectory between the current position and the goal using graph search, considering nodes in the trajectory as sub-goals. The random goals policy chooses a feasible random goal in the map and navigates to the goal with the path planner. The frontier exploration policy implements classical frontier-based exploration [30]: the agent keeps an internal representation of the explored map and computes feasible goals at the frontiers (the limits of the explored map), then it navigates via the path planner to the goal that provides the highest potential information gain (surrounded by the highest number of unknown map cells). We train the CLA policy following the Look Around setup [25].
Instance segmentation. We equip the agent with a Mask2Former model to detect and segment objects in the environment [31]. In particular, we use Mask2Former with a Swin Transformer backbone [32], pre-trained on COCO instance segmentation [33]. For the purposes of this work, we focus on six classes of indoor objects that are commonly found in Gibson and HM3D: couch, potted plant, bed, toilet, tv, table. The size of these categories allows to reduce Mask2Former misdetections and to obtain a sufficiently large resolution of the cropped images to be processed by a captioner. We increase height and width of the bounding box by 10 pixels for each side to avoid cropping the object edges and to include a part of the background (context) in the object crops. After clustering the point cloud into object instances using connected components, we render the RGB images from the collected camera poses in the environment. The bounding boxes are ray-traced from the object clusters in the point cloud to the image plane. To further limit the amount of wrong detections due to the noise in the point cloud or the presence of occlusions, we discard bounding boxes having a confidence lower than $0.7$, an area less than $8000$ pixels, and we filter them using Non-Maximum Suppression with a threshold of $0.8$ in intersection over union [34]. However, due to inaccuracies in the point cloud projection process, such as occlusion, depth estimation, or due to artifacts from the ray-tracing steps, the obtained bounding boxes may not align precisely with the extent of the objects in the image support. Hence, we select only the bounding boxes having intersection over union higher than $0.8$ with the bounding boxes of Mask2Former.
@@
Training and testing images. We launch an exploration episode of 300 steps for each environment in both Gibson and HM3D. Each episode starts from a random position. We equip the agent with CoCa to keep the computational burden of the navigation policy low, as CoCa is 4 times faster than BLIP2. Moreover, we delete repeated samples caused by the agent revisiting certain poses during exploration, to train and test on different images. Since each policy explores the testing environments in a different way, we merge the different sets collected with each policy into a single test set. For each object instance in the test set, we manually annotated a single, concise, and representative caption, by looking at multiple images from different viewpoints. We associate these new annotations with the corresponding object observations collected during agent navigation leveraging the point cloud segmentation of the scene.
Learning captions consistency. We compare two fine-tuning strategies for captioners: standard fine-tuning that minimizes only the captioning loss $\mathcal{L}_{cap}$, and an alternative approach that uses contrastive learning with a triplet loss. We evaluate both approaches against the off-the-shelf models to assess their impact on captioning performance and consistency. We initialize CoCa and BLIP-2 with off-the-shelf weights and we fine-tune them for 10 epochs, applying early stopping if the validation loss does not decrease for 3 consecutive epochs. In the triplet loss, we set the weight $\lambda_{tr}$ to $0.1$, the margin $\epsilon$ to 2, and $d$ as the Euclidean distance. For CoCa, we disable the contrastive loss to prevent penalizing the encoders based on vision-language alignment. Given BLIP-2’s large number of parameters, we apply LoRa adaptation [41] on the query and value projections of the Q-Former module using the Rank-Stabilized version [42]. Training images are augmented with 50% probability using horizontal flip, gaussian noise and affine transformations. A complete list of hyperparameters is provided in Sec. A of the Supplementary Material. Additionally, we compare also with the off-the-shelf version of ChatGPT (o1) [4] and Florence2 [3], fed with the object crops. We use Florence2-base in image captioning mode, and we prompt ChatGPT asking to use no more than 5-7 words, only one adjective per-noun, and we provide some examples of sentences. The full prompt is in Sec. D of the Supplementary Material.

### v2

@@ I Introduction
Visual understanding is a fundamental skill to recognize and describe objects in a scene, enabling robots to navigate and interact with the environment [5]. Despite advances in image captioning [1, 6, 7, 8, 3], a captioner deployed on an autonomous agent often generates wrong or inconsistent descriptions across different views of the same object, especially in the case of occlusions or challenging viewing directions (see Fig. LABEL:fig:main_figure).
Previous works tackled the problem of inconsistent captions from three research directions: agent navigation [9, 10], noisy captions [6, 11, 12], and caption comprehensiveness [13, 14]. Navigation-based captioning methods combine the agent observations at different times during navigation [10]. However, when the agent moves in complex environments, the generated descriptions can be highly inconsistent for the same object. Methods that generate the most appropriate caption for an image either use the alignment between vision and language to select the description from a set of captions [12], or prompt a Large Language Model (LLM) to summarize the captions [13]. Mentioned methods fail in case of inaccurate descriptions that might be selected or propagated to the captions summary.
Compared to previous works, we tackle both the noise presence and the caption coherence as challenges inherently present when a robot navigates a complex indoor environments. To this end, we propose a three-phase framework that combines embodied perception with a self-supervised strategy to retrain the captioner enhancing the performance and increasing the consistency of generated descriptions. In the first phase, the agent equipped with an object detector and a captioner navigates the environment, collecting observed RGB-D images and corresponding object captions. The agent also creates a point cloud representation of the scene and clusters the object instances in 3D, given the output of the object detector. In the second phase, for each 3D segmented object instance, we prompt an LLM using the corresponding captions, their frequency and in-context learning to describe each object in the environment, hence filtering out potentially wrong captions and generating a comprehensive pseudo-label. In the third phase, we specialize the captioner through contrastive learning, maximizing the similarity of representation for different views of the same object while minimizing the similarity with other objects’ representation. The framework design is modular to allow the integration of past and future approaches for exploration, pseudo-caption generation, and environment specialization, fostering fair comparisons at each phase, and offering a foundation to improve caption consistency in challenging environments.
@@
  • A fine-tuning strategy using contrastive learning to improve caption consistency across object views.
@@ II Related Works
Image captioning. Captioners adapt existing models by using additional inputs such as questions about the image, or bounding boxes  [8, 2, 15, 16, 17, 18, 19, 20, 21] or by improving the representations using contrastive learning [1, 22, 7, 8, 3, 6]. To fully exploit the generalization capabilities of foundation models, the majority of methods adjusts the representation of vision-language models fine-tuning tokens or transformer layers, reducing the gap between the pre-trained visual and text representations [2]. For example, X-InstructBLIP and BLIP2 bridge the representation between the image encoder and the text decoder by learning a set of queries to combine visual features with textual features using a lightweight transformer [8, 2]. Other methods like SCA and CAT adapt the SAM architecture [23] through visual chain-of-thoughts [17] or additional cross-attention layers to combine the vision and text representation and predict regional captions [15]. A common approach among captioning methods is to learn the representation through contrastive learning i.e., maximizing the similarity among input modalities such as vision, language, and point cloud [1, 22, 7, 8, 3, 6]. The multimodal alignment of different object representation can be learned either during a pre-training stage enabling downstream tasks such as captioning or classification [7, 8, 3, 6], or during the training to ground the captions to the visual features [1, 22]. Despite the use of contrastive learning on input modalities, methods fail to provide consistent captions across different views of the same object, generating wrong descriptions in the case of challenging viewpoints. In this work, we apply contrastive learning to induce the model to learn close representations of the same object in the latent space while maintaining a margin distance from the representations of other objects, using triplet loss [24].
@@
Embodied captioning. Embodied agents navigate the environment and collect different views of the same object to learn a 3D representation of the scene [9] or to generate a caption [10, 25, 26] . For example, CaBOT decouples the policy that predicts the action to take in the environment, from the captioning model that generates the object description based on the trajectory [10]. Despite generating a single caption for an object, CaBOT has prior knowledge of the best viewpoint in the scene and scenes contain few objects, limiting the application in case of challenging indoor environments like houses or office rooms and in case the annotation of the best viewpoint is not provided. Embodied scene description framework [27] used navigation to find an optimal viewpoint to describe a scene, while Bigazzi et al. [25, 26] minimized overlapping between scene descriptions. However, our objective is to improve the consistency between different views of the same object, instead of improving the variety of description.
@@ III Method
We propose a three-phase pipeline to improve the captioning performance and the consistency of captions across different views of the same object (see Fig. 1). In the first phase, an agent explores an environment following a policy $\pi:S_{t}\rightarrow a_{t}$, where $S_{t}$ is the state and $a_{t}$ the action at time instant $t$, and collects RGB-D observations. The agent is equipped with an instance segmentation model $f_{d}$ and an image captioning model $f_{c}$. $f_{d}$ restricts the RGB observation $I_{t}\in\mathbb{R}^{W\times H\times 3}$ of the agent to detected object instances $o$, $f_{d}:I_{t}\rightarrow\{{l}^{o}_{t},b^{o}_{t},M^{o}_{t}\}$, where $l\in\mathbb{R}^{O}$ is the logits vector of the $O$ classes $f_{d}$ can detect, $b\in\mathbb{R}^{4}$ is the object bounding box, and $M\in[0,1]^{W\times H}$ is the object segmentation mask. $f_{c}$ generates the description of each detected object $f_{c}:I^{\prime o}_{t}\rightarrow c^{o}_{t}$ where $I^{\prime}\in\mathbb{R}^{W^{o}\times H^{o}\times 3}$ is the object crop and $c$ the caption. During exploration, the agent builds a semantic voxel map of the environment consisting of $V$ points, $v_{t}\in\mathbb{R}^{V\times 3}$, and clusters the voxels into object IDs based on the detection logits. In the second phase, an LLM predicts a unique and concise description (pseudo-caption $\tilde{c}$) of each clustered object based on all the corresponding captions collected during exploration. In the third phase, we fine-tune the captioner using pseudo-captions to improve the consistency of captions for the same object instance.
@@ III-A Navigation and point cloud segmentation
Learned policy (Caption-Look Around - CLA). The built voxel map can be used to guide an exploration policy based on maximizing the disagreement between captions of the same object (Fig. 1a). For this reason, we adapt Look-Around [28] policy (inspired by Active Neural Slam [29]) to tackle the captioning task. The policy takes an input $S_{t}$ defined as:
@@
where the first channel of $D^{map}_{t}$ contains a disagreement map with canonical dimension $K\times K$, where $K=128$; the second channel contains the map of the explored environment with the agent’s position superimposed; finally $e_{t}$ contains the global orientation of the agent with respect to the map. For the disagreement map, we use a per-pixel measure of inconsistency between captions, computed as follows: for each object, all captions are encoded via SBERT [30]. Then, the cosine distance among all caption embeddings belonging to the same object is used as a measure of disagreement, which is down-projected onto a 2D image (see Fig. 2d). The policy learns to predict a long-term goal $(x,y)$ in the environment through reinforcement learning using the total disagreement as a curiosity reward [28].
@@ III-B Language-Driven Consistent PSeudocaptioner
The exploration phase assigns a unique ID to the object instances in the built voxel map. However, different captions $c^{o}_{t}$ are still associated with each object instance, and hence they are not consistent across views. We ensure consistency at the label level using a fully self-supervised approach based on prompting an LLM to perform caption distillation. We call our method LD-CPS (Language-Driven Consistent PSeudo-captioner). For each object instance, LD-CPS generates the pseudo-caption $\tilde{c}$ from the list of predicted captions $c_{t}$ as follows: the captions are first pre-processed to remove all uninformative text such as “A picture of…” or “A photo of…” that could be the result of bias in the captioner training or due to the distribution shift between training set and simulated environment. Then, a prompt is devised that asks an LLM to combine the list of all captions, given their frequency i.e., how many times they occur for the same object (Fig. 1d), into a concise caption (Fig. 1e). The prompt also highlights to the LLM the fact that some captions might be inaccurate or wrong. In addition to considering caption frequency, we use in-context learning, by providing a list of examples of how to solve the task inside the prompt. Despite using an LLM to summarize the caption, the proposed approach is different from IC3 [13], since we use the frequency of the captions in the prompt, under the assumption that the noisy data are not the majority. In-context examples are also provided to improve the pseudo-caption distillation task. We show the full prompt in Sec. E of the Supp. Mat.
@@ III-C Learning captions consistency
where $\lambda_{tr}$ is the coefficient weighting the triplet loss.
@@ IV-A Datasets
We validate our framework on two datasets: Gibson [31] and Habitat-Matterport (HM3D) [32]. To avoid using degraded object meshes, we consider a total of 194 scenes with a reconstruction quality between 3 and 5 (the maximum), and we randomly split the scenes into training set (134), validation set (28) and test set (32) keeping the initial distribution of scene quality across the splits. We use the 181 scenes from the Habitat Challenge, split into a training set (126), a validation set (27), and a test set (28). We split data based on the scenes to evaluate the generalization of the framework to environments unseen during training in both Gibson and HM3D.
@@ IV-B Implementation details
Navigation. We import the datasets scenes in the Habitat simulator [33] that allows agent to explore environments and to collect observations. We compare three exploration policies: random goals, frontier exploration and the learned policy (CLA) presented in Sec. III-A. The random goals policy chooses a feasible random goal in the map and navigates to the goal with the path planner. The frontier exploration policy implements classical frontier-based exploration [34]: the agent keeps an internal representation of the explored map and computes feasible goals at the frontiers (the limits of the explored map), then it navigates via the path planner to the goal that provides the highest potential information gain (surrounded by the highest number of unknown map cells). We train CLA following the Look Around setup [28].
Instance segmentation. We equip the agent with a Mask2Former model to detect and segment objects in the environment [35]. In particular, we use Mask2Former with a Swin Transformer backbone [36], pre-trained on COCO instance segmentation [37]. For the purposes of this work, we focus on six classes of indoor objects that are commonly found in Gibson and HM3D: couch, potted plant, bed, toilet, tv, table. The size of these categories allows to reduce Mask2Former misdetections and to obtain a sufficiently large resolution of the cropped images to be processed by a captioner. We increase height and width of the bounding box by 10 pixels for each side to avoid cropping the object edges and to include a part of the background (context) in the object crops. After clustering the point cloud into object instances using connected components, we render the RGB images from the collected camera poses in the environment. The bounding boxes are ray-traced from the object clusters in the point cloud to the image plane. To further limit the amount of wrong detections due to the noise in the point cloud or the presence of occlusions, we discard bounding boxes having a confidence lower than $0.7$, an area less than $8000$ pixels, and we filter them using Non-Maximum Suppression with a threshold of $0.8$ in intersection over union [38]. However, due to inaccuracies in the point cloud projection process, such as occlusion, depth estimation, or due to artifacts from the ray-tracing steps, the obtained bounding boxes may not align precisely with the extent of the objects in the image support. Hence, we repeat the filtering based on IoU 0.8.
@@
Training and testing images. We launch an exploration episode of 300 steps for each environment in both Gibson and HM3D. Each episode starts from a random position. We equip the agent with CoCa to keep the computational burden of the navigation policy low, as CoCa is 4 times faster than BLIP2. Since each policy explores the testing environments in a different way, we merge the different sets collected with each policy into a single test set. For each object instance in the test set, we manually annotated a single, concise, and representative caption, by looking at multiple images from different viewpoints. We associate these new annotations with the corresponding object observations collected during agent navigation leveraging the point cloud segmentation of the scene.
Learning captions consistency. We compare two fine-tuning strategies for captioners: standard fine-tuning that minimizes only the captioning loss $\mathcal{L}_{cap}$, and an alternative approach that uses contrastive learning with a triplet loss. We evaluate both approaches against the off-the-shelf models to assess their impact on captioning performance and consistency. We initialize CoCa and BLIP-2 with off-the-shelf weights and we fine-tune them for 10 epochs, applying early stopping if the validation loss does not decrease for 3 consecutive epochs. In the triplet loss, we set the weight $\lambda_{tr}$ to $0.1$, the margin $\epsilon$ to 2, and $d$ as the Euclidean distance. For CoCa, we disable the contrastive loss to prevent penalizing the encoders based on vision-language alignment. Given BLIP-2’s large number of parameters, we apply LoRa adaptation [45] on the query and value projections of the Q-Former module using the Rank-Stabilized version [46]. A complete list of hyperparameters is provided in Sec. A of the Supp. Mat. Additionally, we compare also with the off-the-shelf version of ChatGPT (o1) [4] and Florence2 [3], fed with the object crops. We use Florence2-base in image captioning mode, and we prompt ChatGPT to use no more than 5-7 words, only one adjective per-noun, and we provide some examples of sentences. The full prompt is in Sec. E of the Supp. Mat.